OpenStudy (nightmarenight):

Solve the equation. If necessary, round your answer to the nearest tenth. 4x2 = 81

OpenStudy (texaschic101):

take the square root of both sides

OpenStudy (texaschic101):

sq rt 4x^2 = sq rt 81 2x = 9 -- now divide both sides by 2 x = 9/2 which equals 4 1/2 or 4.5 any questions ?

OpenStudy (texaschic101):

or if you wanted to....you could divide both sides by 4 giving us : x^2 = 20.25 then take sq rt 20.25 x = 4.5

OpenStudy (nightmarenight):

oooh! thank you c: @texaschic101

OpenStudy (texaschic101):

glad to help :)